OpenStudy (anonymous):

A canoe takes a total of 10 hours to travel from the pier to the boat and back. It travels at a rate of 16 mph and returns at a rate of 64 mph. How far is it from the pier to the boat?

OpenStudy (anonymous):

128 miles

OpenStudy (anonymous):

time = distance/speed for first case t1=x/16....1 and for second t2=x/64....2 here t1+t2=10 so adding 1 and 2 we get 10=x/16+x/64 now solve to get the value of x